OpenStudy (anonymous):

solve for x: 12x-3=3(x+1)+4

OpenStudy (ash2326):

@waltersa Let's work step by step Could you simplify the right hand side? \[12x-3=3(x+1)+4\]

OpenStudy (anonymous):

3x+7

OpenStudy (ash2326):

good, so we have now \[12x-3=3x+7\] Now add 3 to both the sides, what would you get?

OpenStudy (anonymous):

12x=3x+10?

OpenStudy (ash2326):

Great going :) now subtract 3x from both the sides :)

OpenStudy (anonymous):

9x=10

OpenStudy (ash2326):

now divide both the sides by 9 :)

OpenStudy (anonymous):

will it simplify past 10/9

OpenStudy (ash2326):

You could but this is fine

OpenStudy (anonymous):

ok thanks

OpenStudy (ash2326):

no problem :)